OpenStudy (anonymous):

Help me out with a couple problems please? Id really appreciate it :) Link: join.me/marthahearst

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

@ameenbriscoe13 dude you don't know this stuff...

OpenStudy (anonymous):

i do you put x=2 just put 2

OpenStudy (anonymous):

@ameenbriscoe13 ctfu that wasn't right either

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@ageta

OpenStudy (ageta):

what is the question?

OpenStudy (xapproachesinfinity):

what is that link?

OpenStudy (ageta):

i am asking her, can you not interrupt.

OpenStudy (anonymous):

@ageta go to the link and you will see it

OpenStudy (ageta):

i found the link

OpenStudy (ageta):

but it is locked

OpenStudy (xapproachesinfinity):

@ageta i was asking her lol

OpenStudy (ageta):

sorry

OpenStudy (ageta):

9y=x-18 put the equation in the slope-intercept form: y=mx+b 9y/9=x/9-18/9 y=(1/9)x-2 if two lines are perpendicular to each other, they have slopes that are negative reciprocals of each other the slope of the perpendicular line is -9 now we have y=-9x+b using (2,-5), we have -5=-9(2)+b -5=-18+b and b=13 answer: y=-9x+13 using point-slope form, we have y-(-5)=-9(x-2) y+5=-9x+18 y=-9x+13

OpenStudy (xapproachesinfinity):

I don't know what you guys are doing you are randomly doing this!

OpenStudy (xapproachesinfinity):

so here we go the line that was given is 8y=x-16 this one is similar to the other one

OpenStudy (xapproachesinfinity):

again we need to find the slope! can you find the slope

OpenStudy (xapproachesinfinity):

the same steps

OpenStudy (ageta):

9y=x-18 put the equation in the slope-intercept form: y=mx+b 9y/9=x/9-18/9 y=(1/9)x-2 if two lines are perpendicular to each other, they have slopes that are negative reciprocals of each other the slope of the perpendicular line is -9 now we have y=-9x+b using (2,-5), we have -5=-9(2)+b -5=-18+b and b=13 answer: y=-9x+13 using point-slope form, we have y-(-5)=-9(x-2) y+5=-9x+18 y=-9x+13

OpenStudy (xapproachesinfinity):

Did you kick me lol

OpenStudy (xapproachesinfinity):

hey what was the question! i'm trying to help hehe the link does not work or something

OpenStudy (xapproachesinfinity):

@MarthaHearst ??

OpenStudy (xapproachesinfinity):

@ageta same thing happened! i guess the system has some problem

OpenStudy (ageta):

ya it doesn't work.

OpenStudy (ageta):

i just tried it know.

OpenStudy (xapproachesinfinity):

i'm sure the answer is correct

OpenStudy (ageta):

ya me too

OpenStudy (xapproachesinfinity):

oh anyway! the poster wants just answer which is no good in mine opinion!

OpenStudy (ageta):

well i have to sleep because i have to take a test i have life science test 7th grade.

OpenStudy (ageta):

good bye xapproachesinfinity

OpenStudy (xapproachesinfinity):

bye good sleep